AN ADDRESS

IN MEMORY OF

WILLIAM SEYMOUR TYLER,

DELIVERED IN THE

COLLEGE CHURCH AT AMHERST, May 29, 1898,

BY

SAMUEL EDWARD HERRICK, D. D.,

OF MOUNT VERNON CHURCH, BOSTON.

�� ���ν �ληθ�, �� �μν�, �� δ�καια, �� �γν�, �� προ�ιλ�, �� ε�φημα, ε� τισ �ρετ� κα� ε� τισ �παινοσ, τα�τα λογ�ζε�ε.

I approach the pious duty of this hour, my friends and brothers, with feelings of mingled reluctance and alacrity. My heart is in conflict with my judgment. I do not choose this service and yet I cannot decline it. The life and character of Professor Tyler deserve portrayal by a hand no less skillful than his own. The man, the scholar, the teacher, the Christian, all the aspects under which his life may be viewed, demand, for their just and adequate appraisal, a manhood, scholarship, devotion and faith, approximating at least to those of the subject. As we recall the characterizations which he gave to the world of so many of his illustrious contemporaries — Sabin, Dickinson, Stearns, Hitchcock, Humphrey and a host of others — and think how true and just and beautiful these characterizations were, one is tempted to wish that we might have Professor Tyler as sketched by himself. Only that, as Dr. Bartol said of Edwin P. Whipple, “his singular lowliness would have hidden the equally rare splendor of his own gifts.”

But while I may not hope to “apprehend that for which I have been apprehended,” I may claim at least one qualification for the service to which you summon me. Dr. John Brown of Edinburgh said of one of his old teachers, “He had that quality of primary minds of attaching permanently those whom he had relations to. His students never ceased to love him and return to him from all regions of the world. He was in this a solar man, and had his planets pacing faithfully round him.” Professor Tyler was such a solar man. It became my good fortune to fall within the sway of his paternal and gracious influence when I was a boy but fifteen years of age. And thereafter, for more than forty years, he stood to me as the lawyers say “in loco parentis.” For that solar relation of his was not only luminous and genial, but it was far-reaching and permanent. It extended to the very aphelia of his pupils’ careers. Once mutually established, no distance of space and no lapse of time was suffered to rupture it. I can bring you, therefore, an appreciation of our old master. A biography you will not expect, and I shall not pelt you with dates.

Alexander the Great, it is said, held his copy of the Iliad in such honor that he set apart a jewelled casket from his Persian booty to keep it in. I wish I could do something like that for the story of this noble life. But “silver and gold have I none.” I can offer only the simplest tribute of a profound admiration and a life-long affection. I shall not care even if I may seem to any to be partial and prejudiced. I know I shall hear in imagination at many a period of this discourse, the “Pause there!” with which he was wont to arrest us in Homer and Demosthenes and Æschylus and Plato; but I shall not heed even that imaginary injunction of his modesty as I try to tell you Amherst men of the nineties what Professor Tyler was to us of the fifties. Many of you I presume would say that you knew Professor Tyler. And yet you did not know him as we of forty years ago knew him. You knew him as the venerable sage, the ripened saint, the hero who had fought his battles, run his course and kept the faith. We saw him in the heat of the conflict, under the mid-day burden, wrestling with principalities and powers, not always “in heavenly places,” though always with a heavenly purpose and a holy energy.

When a nobly strenuous life comes to its close, it is generally discovered, I think, by those who have witnessed any considerable portion of it, to have had a large element of pathos in it. When it comes to its close, I say: for the glory of the cross is ever posthumous. Only retrospect can give atmosphere, proportion, true tone, to an action or a career. Thirty-five years Boston has waited for St. Gaudens to tell the story of Robert Shaw and his colored troops at Fort Wagner. She has done well. Now it is told fitly and finally. The heroes themselves could not have told, did not know, what they were doing, nor how they were doing it. Heroes never do. Hercules never felt when he was slaying the Nemean lion and plucking the apples of the Hesperides, what the sculptor of the Farnesan statue has put into that pathetic face, and into the serenely patient attitude, with the apples held modestly behind the back, as acquisitions after all of small account. And so Professor Tyler, I think, never realized the pathos or the strenuousness of his own life. It was a life of great hardihood, and mighty courage from the beginning. He gathered his apples, put them behind him, and took the next labor.

He was born amid the simplicities and asperities of Puritan and Colonial life. He has pictured for us the Arcadian scenes of his boyhood, in his semicentennial address given at his birthplace, Harford, Pa., some twenty years ago. To journey back in thought and imagination to the time and place of that boyhood is like going back into some dim, almost legendary region. The little colonies that were going west from New England at the beginning of the century to subdue the forests of Northern Pennsylvania and Central New York, were repeating the earlier experiences of Plymouth and Salem and Haverhill. The wolves were intimately acquainted with their farm yards. Their fires were kindled in their rude stone fire-places from the flint and the steel, or by coals brought from some neighbor’s house half a mile away. It was a time of no carpets on the floors, and no pictures on the walls. The sofa was unknown, save as Cowper had just sung it in “The Task.” Michigan with its furniture factories as yet was not. Kitchen, dining-room, living-room, were one. The good man Tyler, Justice of the Peace, transacted his business, and conducted neighborhood trials in the same apartment where the good wife baked her own pies, and washed her own dishes. Books were few; so few that a new accession to the family library made an epoch. The whole collection would be stored in one of those nondescript pieces of furniture which was book-case, escritoire, and chest of drawers, all in one. The clothing of the household was made from wool raised on the farm, spun in the house, and for the most part by the mother’s own hands. The children of the family milked the cows and drove them to and from the pasture. The one old white horse took the father and mother, and an indefinite number of children upon their infrequent journeys through the obscure wood-paths. And when William came to Amherst in mid-winter of 1829, there was no other way but to pack the boy with his books and clothing, bed and bedding into a sleigh behind the old white horse and drive from Monday morning before day, until Amherst was reached on Saturday at noon.

And yet the strenuousness of this frontier farmer’s life was not without its abundance of cheer. It was softened, moreover, and glorified by dignity, intelligence, fine feeling. “Strength and beauty were in the sanctuary of the Tyler home as they have ever been. Hard work and lofty thought. The parents sought the best things for their children, and taught them to desire the best things for themselves. Their pride and ambition all ran to virtue, integrity, good learning. They belonged to the aristocracy of old New England’s religious and intellectual life. And it would be an interesting speculation, were there time to entertain it, to consider how far and how deeply in the life of our country, spread the influence of the moral and intellectual atmosphere of that one Puritan household. That influence, as it was incarnated in Professor Tyler, — especially as it mellowed, and softened, and grew gracious through the years when his own children were growing up around him — made itself felt upon generations of plastic students, who carried its undying impress to all lands under the sun.

In the middle of the century Amherst was poor. There were few students here from wealthy families — one or two perhaps upon every row of benches in the old Greek room — just enough to suggest to the rest of us what a pleasant and delightful thing it must be, to always have money in the pocket, and never have to wear threadbare clothes. But there sat the teacher, who, as we all knew, thought never of the clothes, but only and always of the man. We all knew just what things were at a premium in that classroom, and just what coin would not be current. He dearly loved neatness, accuracy, adequacy in a boy’s rendering, but he loved more an honest endeavor, even though it were accompanied with dullness in the intellectuals. The moral element was uppermost in all his judgments. There was what one of his old pupils called “a genial intolerance of slovenly rendering,” when he knew that there was a good intent. But when there was any shamming, or any unreality, the intolerance was no longer genial, but satirical and withering. There was no deceiving him, nor did he allow us to deceive ourselves. We knew at the end of a recitation just what marks would go upon the faculty records, and we knew that they would be righteous; righteous and kind. I remember once when we were reading the Iliad, one of our number had translated a passage with a fluency, and at the same time with an adequacy which in his case were altogether unaccountable, save upon the supposition that he was reading at sight from an inserted page of his “Bohn.” Things went swimmingly until he came to the last sentence, which began with the Greek name of Juno, simply transferred, instead of being translated. Sublimely audacious, he read on, “Here the goddess interposed,” etc. “Pause there!” said the Professor. “He-re, the goddess interposed, isn’t it? Next!” That glib translation, like Achilles, was vulnerable only in its heel; but no need to further puncture it with question or unmask it by comment. It was all as kindly as a flash of white light, and as thorough. There was the swift, electric illumination; the immediate and automatic condemnation of a day of judgment, without the possibility of extenuation or appeal.

And yet his severity was much like the coat of a russet apple. It was well understood that beneath the austerity of his seeming there was an immense fund of paternal feeling that could be relied upon. He “dealt justly, and loved mercy.” He was humane to a degree. He loved to find some point of view from which peccadillos might be regarded as excusable. He was a favorite mediator therefore upon those occasions of conflict which would now and then arise in the college community, between the conduct of the students and the judgment of the faculty. If he held a brief from us youngsters, we felt sure that the case would go well at Court; if he refused to be retained by us, we were equally certain that the decision of the Court would be righteousness tempered by his leniency.

As a teacher, he brought his own personality to bear upon his pupils in a very remarkable way. The university method insulating to all personality had not so much as knocked at the door of his class-room. What he taught was all unconsciously and yet profoundly saturated with the spirit of the man himself. First and foremost a class-room exercise was with him an act of religion. Professor Agassiz once said, “I never make the preparations for penetrating into some small province of Nature hitherto undiscovered, without breathing a prayer to the Being who hides His secrets from me only to allure me graciously on to the unfolding of them.” Professor Tyler might have said the same concerning his daily excursions with his class into the domain of Greek Literature and Philology. There was a certain reverent humility in his entering upon the exercise of the hour, as of one who is entering a temple to worship. As Charles Lamb would have had “a grace before Milton, a grace before Shakspeare, a devotional exercise proper to be said before reading the Fairy Queen,” so Professor Tyler approached a reading from the Iliad, the Phaedo, or the Prometheus Bound. He was not more earnest, or more reverent in reading the Hebrew prophets and the Christian Apostles on Sunday than he was in interpreting Æschylus or Plato on Monday. With him these also were prophets and apostles, though lacking it might be some higher and finer degree of inspiration. Their teachings were true theological preliminaries — threshold matters. With them he felt himself to be treading the outer and more ancient precincts of Christian Theology. In their moral and religious teachings, he saw the projecting lines of a divine intent. These were not merely the “unconscious prophecies of heathendom,” but Jehovah was there, “laying the beams of His chambers in the waters.” If we may not say that in those early days he had actually seized, yet I am persuaded that he was already glimpsing, that large and profound conception of the religious development of the race, of the God-ward struggle of the human soul, by which the scientific study of comparative religion has in these later times thrown into solution the old theological systems preparatory to some new crystallization. He made us feel that these old Greeks were vitally a part of our world of to-day — not only that they were endowed with the same human nature as ourselves — but that as the growth of a thousand years ago still abides in the substance and beauty of some venerable oak yet vigorous, so their speech still vitalizes ours; their wisdom and experience are in our institutions and philosophies; and even their ethical and religious thinking must be taken account of to-day by any who would understand our own. He made us feel that Greek History and Greek Literature were a most precious and vital contribution to the providential education of the world, and that Plato and Æschylus, no less than Moses and David, were “schoolmasters unto Christ.” To be in his class-room and to catch his thought was to be liberated and liberalized, and made expectant and ready for all progress. He knew how to collate a page of Demosthenes with a column of yesterday’s newspaper; Socrates with Horace Greeley. I take it that not much is known to-day of the finical way of studying Greek which generally prevailed in the college class-rooms of forty years ago. That is largely thrust back into the preparatory school, as it should be. But at that time it was necessary for the professor’s desk to be more or less of a dissecting table. A good deal of time had to be taken up in the minute analysis of orthographic forms. It was a “house of bondage,” both for the teacher and the pupil. But notwithstanding this, Professor Tyler always managed to deliver us from the thraldom of the letter into the freedom of the spirit. How his rugged face would light up as if the teacher, the class, the whole hour, were redeemed when some bright pupil, like Mather or Ward, seized upon the spirit of a sentence of the Gorgias, and rendered it with idiomatic freedom into a bit of contemporaneous slang, or into a phrase from some recent speech of Wendell Phillips, or of Rufus Choate. And so he made us feel that the old world was still the living world, and that we were not so much walking and talking with the Greeks, as that the Greeks were walking and talking with us. His course with any single class was in itself a generous curriculum. He taught us rhetoric with Demosthenes, philosophy with Plato, logic and ethics with Socrates, literature with Homer, theology with Æschylus and Euripides and Plutarch.

Since these words were written I have been favored with the sight of an autobiographical sketch of his life, left in MS. by Professor Tyler, which, in a single paragraph, so illustrates this spirit and temper of his class-room work that I must quote it. “In all the changes of teachers and the times for more than half a century; in required and optional studies; in regular lessons and in reading at sight; by recitations and lectures; in the text-books which I have edited, and the Socratic conversations by question and answer’ which I have habitually held with my classes, my end and aim has been one and the same, — not to teach words only, but words in their inseparable connection with things, and thoughts, I take it, are the greatest and best things; not to teach the lesson only, or the language only, or the literature only, or the life of the Greeks only, but the lesson and the language, and the literature, and the life; and that not of the Greeks only, but of mankind as illustrated by that of the Greeks; not only that I might make Grecians but scholars, and not scholars but men, and not only men but Christians, for the Christian is the highest style of man.’ So may Greek always be taught in Amherst College.” To which prayer of our dear old master, I am sure, every true son of Amherst will fervently say Amen!

As a teacher Professor Tyler was thus a true prophet — a John Baptist — the herald and forerunner of a better day for Greek learning; preparing the way for, and making possible, Mather and his broadening work. We may be thankful that he was permitted to see so much of what he longed for, — the superstructure whose foundations he toiled so strenuously, so humbly, so magnanimously to lay. He saw it without jealousy and with great delight. The thought, “He must increase, but I must decrease,” brought nothing but joy to the humble old scholar’s heart. In this last third of the century, speaking roughly, an era of great expansion in all directions, a period of great fortunes, great industrial corporations, great mercantile trusts, great gifts to causes of religion and literature, great institutional churches, great universities, great libraries, and perhaps we may add, great learning — under the stimulus of specialization — it is easy to forget or undervalue the prophet who “stood crying in the wilderness, with garment of camel’s hair, and leathern girdle about his loins, feeding upon locusts and wild honey.” But whatever gospel of more generous culture may have dawned, or may yet dawn, upon Amherst College, she can never forget the one life which more than any other made such a new day possible. The time will come, if indeed it has not already come, when Amherst will be said to have had an Heroic Age; when men will point to the megaliths of her foundations, and discuss the stature and character of her early builders. What name will then be likely to have gathered to itself more of high renown than that of our great Greek teacher through well-nigh six decades? Then will be seen as we cannot see it, something of the power, the use, the beauty, the pathos of this heroic life. Many of his specific labors will doubtless be forgotten. This or that great blow by which he wrought may not be in evidence. His classical editions may become obsolete. The learning — marvellous for his day, — with which he entered into the great fight over the Homeric question, and beneath which he buried, as it seemed forever, the paste-pot and scissors theory, and satisfied us all that “Homer wrote the Iliad, the whole Iliad, and the Iliad as a whole; Homer wrote the Odyssey, the whole Odyssey, and the Odyssey as a whole,” will likely be lost sight of in the vast advances of criticism and philology. But that huge stretch of foundation work — binding into itself four complete administrations of college history — that work done on a meager salary — always for love and not for pay — work done under limitation — far from libraries — with great difficulties and almost no facilities — often thankless, but always patient, generous, faithful — that work, I say, will be seen in coming time, as the most momentous fact of our Alma Mater’s first century.

Professor Tyler may not have been the most conspicuous figure of any particular decade, but he has been by all means, taking the life of the college as a whole thus far, its most essential factor, the commune vinculum which has bound that life through all its administrations and eras into unity; I fear he would hardly forgive me for saying which has bound its separate hero-songs into a single epic.

There can be no fair estimate of his work here which does not take account of Professor Tyler in the College pulpit. As I have already intimated, he was greatly in demand as a preacher of occasional and memorial discourses. His protracted and intimate relations with great educational interests, especially in Central Massachusetts, his personal acquaintance with those who founded, patronized and filled the chairs of superintendence and instruction, and especially his insight into character, and his rare skill in presenting, not an outline nor a profile, but an all-round, stereoscopic, and at the same time an interior and sympathetic, view of his subject, made him the favorite literary biographer of the whole region. He was known to have great store of myrrh and nard, and amber and byssus, and so he became a sort of embalmer-general for the schools and churches of Franklin and Hampshire and Hampden counties. The bulk of his published discourses are probably of this character. But these were not the discourses which revealed his power. We remember him not as an orator, but as a prophet of God, and a preacher of righteousness. He was to Amherst, and about the same time, what the great President James Walker was to Harvard. In their literary quality his sermons were fine and strong; of great compactness and without redundancy; clear and sharp-lined; often epigrammatic, with shot-like sentences whose impact was felt and remembered. But their chief charm and power lay not so much in their distinctly literary and scholarly character, as in their profoundly ethical and spiritual force. He would seize upon some great fundamental principle in morals — he would take some short and pithy text, like “Let integrity and uprightness preserve me,” and flash it like an illuminating search-light all around the horizon, penetrating with it successively the realms of individual, social and national life, until every soul within the hearing of his voice knew what was the evil of the time, and what its anti-septic and its cure.

And his presentation of his theme was characteristic. It was synthetic and cumulative. His conception did not grow up before us as by some self-moved process from within — but it was built up; wrought, not cast; now by touches as graceful as though he were a Benvenuto Cellini carving a hanap, and now by blow on blow like Siegfried forging his sword. When he had finished, the result was a work, massive, concrete, full of “strength and beauty.” Often indeed, it was what the old Hebrew prophets, — whom he closely resembled in spirit and method — called “a burden.” A burden carried first in his own soul, and then by this simple, earnest, cumulative process of disclosure, — revealing its righteousness and its reasonableness — it became a burden rolled upon the souls of his hearers. So he used that day of judgment — the great financial crisis of ’57 — as an occasion for settling his pupils in fundamental righteousness. So he used those days that quickly followed, when rebellion was being incubated at the South, and men at the North were divided and wavering, to hold Amherst College to loyal and patriotic integrity, and prepare her for the loving sacrifice which was so soon to be demanded of her — which was so soon to be demanded of himself, in the offering of his own first-born. And so, in a sense I suppose hardly contemplated by the author of the Epistle to the Hebrews, in his use of the phrase, our Professor, from the College pulpit “wrought righteousness” and “turned armies to flight.” His words, robust and brawny, did make righteous men, and fashion patriots.

But quite as patent in his pulpit-work as the spirit of the Hebrew prophet, was that of the Christian Apostle. The history of Amherst College has been marked throughout by frequently recurring periods of deep religious feeling. Such periods he looked for. He recognized what James Martineau calls “the tides of the Spirit.” He was keen to detect the signs of their approach, and when they came he was in full readiness for them. It can hardly be said that he was first of all to be seen riding upon their crest. He was no drift-wood, sporting on the surface-motion of the hour. His nature was too deep and strong to curvet and caracole upon the eddies of emotional excitation. Homer and Plato must hold their steadfast place in the class-room through the week. He did not believe that men were to be saved by neglecting their appointed work. There was, it may be, some touch of tenderer light upon the monotony of the curriculum, some gentler and quite indescribable feeling suffused “the dull dream of sense and custom.” But when the Sunday came and it was the Professor’s turn to preach, it was as though our John the Baptist had become St. John the Apostle. Not that his grand customary ethic was laid aside or forgotten, but the moral motive was suffused by something distinctively more personal and affectionate. “That which was from the beginning, which we have heard, which we have seen with our eyes, which we have looked upon, and our hands have handled of the Word of life; that which we have seen and heard declare we unto you, that ye also may have fellowship with us; and truly our fellowship is with the Father, and with His Son, Jesus Christ.” There was the same synthetic, cumulative, method, the same upbuilding of his theme by tender touches and by mighty strokes on this side and that, until his sermon had become like a bed of his own Pennsylvania anthracite thoroughly on fire, and his conception was all before us, glowing with the steady ardor of the love of Christ, and we were made to realize that the Christian motive transcended in its grandeur all other motives to scholarship, to manhood, and to all virtue.

But great as Professor Tyler was in the functions of teacher and preacher, he was greatest of all as the man and the friend. I have often thought, since I left College, that the feeling of many of his pupils while under him erected an imaginary barrier between him and themselves, which, unreal as it was, neither could quite break through. He certainly did his best to annihilate all distance and all fear. But it was when the pupil had gained his independent footing, and was doing his own work in the world, and the eyes of each looked level into the eyes of the other, and each greeted the other as a fellow-worker in the kingdom of God on earth, that we his old pupils realized what he was and what he always had been to us. Distance sometimes does more than lend enchantment. It rectifies our judgments. By some inexplicable paradox it discloses sympathies, and reveals in their true significance relations that have been but half apprehended and imperfectly utilized. Professor Tyler, great as he had been in the Greek room and in the chapel pulpit, became still more to his pupils, as they thought of him from the banks of the Oregon and the Tigris, by the Golden Gate and the Golden Horn, in the cities of the East and on the prairies of the West, in the camp and in the kraal, in their pulpits and professors’ chairs, in their editorial sanctums and courts of law. He has been the fixed star to which our eyes have turned from whatsoever quarter of the world, — the star of perpetual apparition in the old Amherst sky. To think of Amherst, has been, first and foremost, to think of him. The picture of “the elder brother” in our Lord’s parable interprets his life, not by its likeness, but by its contrast. He has been our elder brother who has staid patiently at home, working the old farm, training the younger children, the stay and comfort of the venerable parent, the curator of our family history, keeping up the home feeling and the home traditions, and when we prodigals have returned from our wanderings, who so ready as he to kill the fatted calf and make merry for our home-coming? Never once thinking of any poor kid which had been denied him, or of all that he had denied himself, but, with all his heart, rejoicing in the prosperity, or sorrowing in the adversity of his far-roving brethren, — true imitator and counterpart of the great Elder Brother of us all!

Venerable and beloved Master! Thou hast taught us by thy self-denial and patience; by thy courage and humility; by thy faithfulness and true manhood: — “Whatsoever things are true; whatsoever things are honorable; whatsoever things are just; whatsoever things are pure; whatsoever things are lovely; whatsoever things are of good report, — if there be any virtue, and if there be any praise” — τ� καλ� κα� τ� �γαθ� — these things thou hast taught us — in all fidelity to thy Master and ours!